OpenStudy (anonymous):

what do i do to get x = 0 or x = 9 from here x2 + 9x = 0

OpenStudy (anonymous):

Factor out a common term.

OpenStudy (anonymous):

question

OpenStudy (anonymous):

?

OpenStudy (anonymous):

is it (x - 9)(x + 9)

OpenStudy (anonymous):

No, if you FOIL those terms you will get \(x^2-81\). There is a single factor that is common to both terms. \(x^2+9x\)

OpenStudy (anonymous):

i'm confused

OpenStudy (anonymous):

What is \(x \times x\)?

OpenStudy (anonymous):

x^2

OpenStudy (anonymous):

What is \(9 \times x\)?

OpenStudy (anonymous):

so is it (x + 3)(x - 3)

OpenStudy (anonymous):

Nope. If you FOIL those factors you will get \(x^2-9\). Not quite what you are looking for. You may be over thinking it. What is \(9\times x\)?

OpenStudy (anonymous):

(x + 3)^2

OpenStudy (anonymous):

You are jumping ahead. Answer my questions and you will see a pattern.... \(x\times x=x^2\) What is \(9 \times x\)?

OpenStudy (anonymous):

9x

OpenStudy (anonymous):

Great! So... \(x\times x = x^2\) \(9 \times x = 9x\) The factors are on the left and the product is on the right. What factor does each term have?

OpenStudy (anonymous):

factor \(\times \) factor = product

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

You're welcome.